OpenStudy (anonymous):

Two forces with magnitudes of 25 and 30 pounds act on an object at angles of 10° and 100° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (amistre64):



OpenStudy (amistre64):

might want to divvy up the angles into xy parts such that: x = rcos(t) and y = rsin(t)

OpenStudy (amistre64):

personally, i would negate the 100 by 30 vector and add it to the other one

OpenStudy (amistre64):



OpenStudy (anonymous):

i have still no clue how to solve this problem..

OpenStudy (amistre64):

:) we could go a longer route; find the up and over forces and compare them for the 30 by 100 vector, x = 30 cos(100). y = 30sin(100) for the 25 by 10 vector, x = 25 cos(10). y = 25sin(10) the net values of the x and y parts will equate to the resultant vector components and force

OpenStudy (amistre64):

so resultant is equal to the sum of the forces in the given x and y directions x = 30 cos(100). y = 30sin(100) x = 25 cos(10). y = 25sin(10) Rx = 30cos(100) + 25cos(100) Ry = 30sin(10) + 25sin(10)

OpenStudy (amistre64):

the angle would then be the arctangent of Rx/Ry

OpenStudy (amistre64):

the angle would then be the arctangent of Ry/Rx

OpenStudy (amistre64):

and the force is the "length" of the vector: F = sqrt(Rx^2+Ry^2)

OpenStudy (anonymous):

so when i find F i get my answer?

OpenStudy (amistre64):

Find the direction AND magnitude arctan(Ry/Rx) sqrt(Rx^2+Ry^2)

OpenStudy (anonymous):

okat thank you veryy much! :)